A square coil of side l=10cm is mounted on a fixed insulating axis passing through the midpoints of a pair of opposite sides . A very long straight conductor parallel to the axis and at a distance of d=12cm from it carries a current of 65A. When the coil is in the positionin which its opposite sides parallel to the axis are equidistant from the long straight conductor , Find the torque acting on it if a current of 78A passes through it?

Grade:12

1 Answers

ROSHAN MUJEEB
3 years ago
Length of a side of the square coil, l = 10 cm = 0.1 m

Current flowing in the coil, I = 12 A

Number of turns on the coil, n = 20

Angle made by the plane of the coil with magnetic field, θ = 30°

Strength of magnetic field, B = 0.80 T

Magnitude of the magnetic torque experienced by the coil in the magnetic field is given by the relation,

T = n BIA sinθ

Where,

A = Area of the square coil

l × l = 0.1 × 0.1 = 0.01 m2

∴ T = 20 × 0.8 × 12 × 0.01 × sin30°

= 0.96 N m

Hence, the magnitude of the torque experienced by the coil is 0.96 N m.